# Supplement

### v3

@@ 6 Discussion
$[w_{\mu}|{\rm others}]\sim MVN\{\left(\frac{N\mathbf{B}^{t}\mathbf{B}}{\sigma^{2}}+h_{\mu}\mathbf{P}_{\alpha}\right)^{-1}\frac{1}{\sigma^{2}}\sum_{i=1}^{N}\mathbf{B}^{t}D_{i},\left(\frac{N\mathbf{B}^{t}\mathbf{B}}{\sigma^{2}}+h_{\mu}\mathbf{P}_{\alpha}\right)^{-1}\}$
@@
$\propto\exp(-\frac{||Y_{i}-\mathbf{B}(w_{\mu}+\sum_{p=1}^{K}\xi_{ip}\psi_{p})||^{2}}{2\sigma^{2}}-\frac{\xi_{ik}^{2}}{2\lambda_{k}})$
$\propto\exp(-\frac{||Y_{i}-\mathbf{B}(w_{\mu}+\sum_{p\neq k}^{K}\xi_{ip}\psi_{p})-\mathbf{B}\psi_{k}\xi_{ik}||^{2}}{2\sigma^{2}}-\frac{\xi_{ik}^{2}}{2\lambda_{k}})$
@@
$f(\xi_{ik}|{\rm others})$ $\propto\exp(-\frac{||P_{ik}-\mathbf{B}\psi_{k}\xi_{ik}||^{2}}{2\sigma^{2}}-\frac{\xi_{ik}^{2}}{2\lambda_{k}})$
$\propto\exp(-\frac{||\mathbf{B}\psi_{k}||^{2}\xi_{ik}^{2}-2P_{ik}^{t}\mathbf{B}\psi_{k}\xi_{ik}}{2\sigma^{2}}-\frac{\xi_{ik}^{2}}{2\lambda_{k}})$
$\propto\exp[-\frac{1}{2}\left\{\xi_{ik}^{2}\left(\frac{||\mathbf{B}\psi_{k}||^{2}}{\sigma^{2}}+\frac{1}{\lambda_{k}}\right)-2\frac{P_{ik}^{t}\mathbf{B}\psi_{k}}{\sigma^{2}}\xi_{ik}\right\}]$
@@
$[\xi_{ik}|{\rm others}]\sim N\{\left(\frac{||\mathbf{B}\psi_{k}||^{2}}{\sigma^{2}}+\frac{1}{\lambda_{k}}\right)^{-1}\frac{P_{ik}^{t}\mathbf{B}\psi_{k}}{\sigma^{2}},\left(\frac{||\mathbf{B}\psi_{k}||^{2}}{\sigma^{2}}+\frac{1}{\lambda_{k}}\right)^{-1}\}$
@@
$\propto\exp(-\frac{1}{2\sigma^{2}}\sum_{i=1}^{N}||Y_{i}-\mathbf{B}(w_{\mu}+\sum_{k=1}^{K}\xi_{ik}\psi_{k})||^{2})\times\exp(-\frac{1}{2}\sum_{k=1}^{K}h_{k}\psi_{k}^{t}\mathbf{P}_{\alpha}\psi_{k})\times\mathbbm{1}(\boldsymbol{\Psi}\in\mathcal{V}_{K,Q})$
@@
$f(\boldsymbol{\Psi}|{\rm others})$ $\propto\exp\left\{-\frac{1}{2\sigma^{2}}\sum_{i=1}^{N}||R_{i}-\mathbf{B}\boldsymbol{\Psi}\xi_{i}||^{2}-\frac{1}{2}\operatorname{tr}(\mathbf{H}\boldsymbol{\Psi}^{t}\mathbf{P}_{\alpha}\boldsymbol{\Psi})\right\}\times\mathbbm{1}(\boldsymbol{\Psi}\in\mathcal{V}_{K,Q})$
$\propto\exp\left\{-\frac{1}{2\sigma^{2}}\sum_{i=1}^{N}\left(||R_{i}||^{2}-2R_{i}^{t}\mathbf{B}\boldsymbol{\Psi}\xi_{i}+||\mathbf{B}\boldsymbol{\Psi}\xi_{i}||^{2}\right)-\frac{1}{2}\operatorname{tr}(\mathbf{H}\boldsymbol{\Psi}^{t}\mathbf{P}_{\alpha}\boldsymbol{\Psi})\right\}\times\mathbbm{1}(\boldsymbol{\Psi}\in\mathcal{V}_{K,Q})$
$\propto\exp\left\{-\frac{1}{2\sigma^{2}}\sum_{i=1}^{N}\left(-2R_{i}^{t}\mathbf{B}\boldsymbol{\Psi}\xi_{i}+||\mathbf{B}\boldsymbol{\Psi}\xi_{i}||^{2}\right)-\frac{1}{2}\operatorname{tr}(\mathbf{H}\boldsymbol{\Psi}^{t}\mathbf{P}_{\alpha}\boldsymbol{\Psi})\right\}\times\mathbbm{1}(\boldsymbol{\Psi}\in\mathcal{V}_{K,Q})$
$\propto\exp\left\{-\frac{1}{2\sigma^{2}}\left(-2\sum_{i=1}^{N}R_{i}^{t}\mathbf{B}\boldsymbol{\Psi}\xi_{i}+\sum_{i=1}^{N}||\mathbf{B}\boldsymbol{\Psi}\xi_{i}||^{2}\right)-\frac{1}{2}\operatorname{tr}(\mathbf{H}\boldsymbol{\Psi}^{t}\mathbf{P}_{\alpha}\boldsymbol{\Psi})\right\}\times\mathbbm{1}(\boldsymbol{\Psi}\in\mathcal{V}_{K,Q})$
We now introduce more notation: score matrix $\Xi\in\mathbb{R}^{N\times K}$, where each row is the score vector $\xi_{i}$, and residual matrix $\mathbf{R}\in\mathbb{R}^{N\times M}$, where each row is the residual vector $R_{i}$.
$f(\boldsymbol{\Psi}|{\rm others})$ $\propto\exp\left[-\frac{1}{2\sigma^{2}}\left\{-2\operatorname{tr}(\mathbf{R}^{t}\mathbf{B}\boldsymbol{\Psi}\Xi)+\operatorname{tr}(\Xi^{t}\boldsymbol{\Psi}^{t}\mathbf{B}^{t}\mathbf{B}\boldsymbol{\Psi}\Xi)\right\}-\frac{1}{2}\operatorname{tr}\{\mathbf{H}\boldsymbol{\Psi}^{t}\mathbf{P}_{\alpha}\boldsymbol{\Psi}\}\right]\times\mathbbm{1}(\boldsymbol{\Psi}\in\mathcal{V}_{K,Q})$
$\propto\exp\left\{\frac{1}{\sigma^{2}}\operatorname{tr}(\mathbf{R}^{t}\mathbf{B}\boldsymbol{\Psi}\Xi)-\frac{1}{2\sigma^{2}}\operatorname{tr}(\Xi^{t}\boldsymbol{\Psi}^{t}\mathbf{B}^{t}\mathbf{B}\boldsymbol{\Psi}\Xi)-\frac{1}{2}\operatorname{tr}(\mathbf{B}\boldsymbol{\Psi}^{t}\mathbf{P}_{\alpha}\boldsymbol{\Psi})\right\}\times\mathbbm{1}(\boldsymbol{\Psi}\in\mathcal{V}_{K,Q})$
$\propto\exp\left\{\frac{1}{\sigma^{2}}\operatorname{tr}(\Xi\mathbf{R}^{t}\mathbf{B}\boldsymbol{\Psi})-\frac{1}{2\sigma^{2}}\operatorname{tr}(\Xi\Xi^{t}\boldsymbol{\Psi}^{t}\mathbf{B}^{t}\mathbf{B}\boldsymbol{\Psi})-\frac{1}{2}\operatorname{tr}(\mathbf{B}\boldsymbol{\Psi}^{t}\mathbf{P}_{\alpha}\boldsymbol{\Psi})\right\}\times\mathbbm{1}(\boldsymbol{\Psi}\in\mathcal{V}_{K,Q})$
$\propto\exp\left\{\operatorname{tr}\left(\frac{\Xi\mathbf{R}^{t}\mathbf{B}\boldsymbol{\Psi}}{\sigma^{2}}-\frac{\Xi\Xi^{t}\boldsymbol{\Psi}^{t}\mathbf{B}^{t}\mathbf{B}\boldsymbol{\Psi}}{2\sigma^{2}}-\frac{\mathbf{B}\boldsymbol{\Psi}^{t}\mathbf{P}_{\alpha}\boldsymbol{\Psi}}{2}\right)\right\}\times\mathbbm{1}(\boldsymbol{\Psi}\in\mathcal{V}_{K,Q})$
The final line above is the form found in Result 1, which is not of a known distributional family.
S3 Simulation results
S3.1 Background
@@
S3.2 Fixed effects comparison
We compare the fixed effects estimates of $\mu(t)$ produced by FAST to those from the 3 comparator methods. We compare ISE of the posterior estimate and aggregate coverage of equal-tailed $95\%$ credible intervals, each calculated using the same procedures outlined for the FPCs in Section 4.
@@
Figure S3: Kernel smoother of coverage probabilities of $95$% credible intervals of the true mean $\mu(t)$ for FAST and GFSR. Row indicates simulation scenario, S1 then S2. Distribution means: horizontal solid lines; nominal $95$% level: horizontal dotted lines.
@@
Figure S4 visualizes the distribution of score MSE over simulations. FAST is has consistently similar or lower MSE when compared to the existing methods. GFSR has much larger MSE for those scores associated with $\phi_{1}(t)$ in S1. The order of this MSE is close to $19600140^{2}=$ 19600, supporting our previous assertion that the mean $\mu(t)$ is being projected onto $\phi_{1}(t)$ and $\phi_{3}(t)$ by this implementation.
Figure S5: Kernel smoother of coverage probabilities of $95$% credible intervals of the true scores, $\xi_{ik}^{b}$, for FAST, GFSR, POLAR, and VMP. First row: S1; second row: S2. Each column: corresponding eigenfunction. Distribution means: horizontal solid lines; nominal $95$% level: horizontal dotted lines.
Figure S5 displays the kernel smooth of the vectors of coverage proportions for FAST and all 3 comparator methods by FPC (columns) and simulation scenario (rows). FAST is the only implementation which produces consistently close to nominal coverage for all FPCs under both scenarios.
S3.4 Multilevel extension
@@
$\phi_{k}^{(1)}(t)$ =$=\{\sqrt{2}\sin(2\pi t),\sqrt{2}\cos(2\pi t),\sqrt{2}\sin(4\pi t),\sqrt{2}\cos(4\pi t)\};\quad\lambda^{(1)}_{k}=0.5^{k-1}$
$\phi_{l}^{(2)}(t)$ =$=\{1,\sqrt{3}(2t-1),\sqrt{5}(6t^{2}-6t+1),\sqrt{7}(20t^{3}-30t^{2}+12t-1)\};\quad\lambda^{(2)}_{l}=0.5^{l-1}$
@@
Figure S7: Kernel smoother of coverage probabilities of $95$% credible intervals of the true FPCs for FAST and GFSR from the multilevel simulations. Row indicates hierarchical level, while columns correspond to FPCs. The horizontal dotted lines indicate the nominal $95$% level.
@@
Figure S8: Kernel smoother of coverage probabilities of $95$% credible intervals of the true scores, $\xi_{ik}^{b},\zeta_{ijk}^{b}$, for FAST and GFSR from the multilevel simulations. Row indicates hierarchical level, while columns correspond to FPCs. The horizontal dotted lines indicate the nominal $95$% level.
S4 Simulation sensitivity analyses
@@
S4.1 Chosen number of FPCs $K$
@@
Figure S9 indicates setting $K=2$ produces biased estimates of $\phi_{1}(t),\phi_{2}(t)$ as expected, though it also shows that increasing $K\geq 3$ does not result in substantial bias estimating the first 3 FPCs.
@@
Figure S10 similarly shows that $K=2$ produces non-optimal inference for $\phi_{1}(t),\phi_{2}(t)$, with inferences for the first 3 FPCs unaffected when $K>3$.
S4.2 Chosen spline dimension $Q$
@@
S4.3 Chosen penalty parameter $\alpha$
We finally vary penalty hyperparameter $\alpha$, which dictates the proportion of the smoothing penalty used by FAST which focuses on absolute value rather than second-order variation. We vary between $\alpha=0.01$ and $\alpha=0.3$, keeping $\alpha<0.5$ to ensure the majority of the penalty is focused on the ”wiggliness” which should be the core of any smoothing penalty. We evaluate FAST under each $\alpha$ value using ISE and coverage of the FPCs as described in Section 4.
@@
S4.4 Timing by $Q,K$
Given that it appears that estimation accuracy and inference are consistent when $K$ and $Q$ are sufficiently large, we evaluate the effect of increasing these quantities on the computation time of FAST. For this experiment, we focus on S1, the more computationally complex simulation scenario based upon CGM data. For each combination of $K\in\{3,4,5\}$ and $Q\in\{20,30,40\}$, we timed FAST on the same personal laptop described in Section 4 (2023 MacBook Pro with Apple M2 Max@@@@3.49 GHz and 32GB of memory).
@@
From Table S1, we find that FAST does require additional computation time when $Q,K$ are increased, particular when $K$ is made larger. However, the change in computation time is relatively marginal, and FAST is still more efficient in every case than the comparator MCMC-based methods tested on the same problem in Table 1.
S5 Additional CGM Analyses/Details
S5.1 Variability explained - Single Level
@@
S5.2 Bayesian FPCA of randomly-chosen CGM within diets
@@
Figure S16: Bayesian FPCA results for the first three PCs (each column corresponds to one FPC) for each of the four diets (each row corresponds to one diet). X-axis: time from the start of the meal. Black curves: posterior mean; red curves: three samples from the posterior of the PCs; shaded areas: pointwise $95$% credible intervals.
@@
Figure S17: Eigenvalue and percent-variability estimates from the Bayesian FPCA models fit to each of the four diets (each line and color corresponds to one diet). X-axis: eigenfunction corresponding to eigenvalue/percent variance explained. All estimates are presented with their corresponding $95$% credible intervals.
As for the aggregate data, we fit each model using $5000$ iterations with the first $2500$ discarded as burn-in. Employing the routine from Section 2.5 to assess convergence of the scores and FPCs, we found final Gelman-Rubin statistics were $\leq 1.02$.

### v4

@@ 6 Discussion
$w_{\mu}|{\rm others}]\sim MVN\{\left(\frac{N\mathbf{B}^{t}\mathbf{B}}{\sigma^{2}}+h_{\mu}\mathbf{P}_{\alpha}\right)^{-1}\frac{1}{\sigma^{2}}\sum_{i=1}^{N}\mathbf{B}^{t}D_{i},\left(\frac{N\mathbf{B}^{t}\mathbf{B}}{\sigma^{2}}+h_{\mu}\mathbf{P}_{\alpha}\right)^{-1}\}$
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
We now introduce more notation: score matrix $\Xi\in\mathbb{R}^{K\times N}$, where each row is the score vector $\xi_{i}$, and residual matrix $\mathbf{R}\in\mathbb{R}^{N\times M}$, where each row is the residual vector $R_{i}^{t}$.
$f(\boldsymbol{\Psi}|{\rm others})$ $\propto\exp\left[-\frac{1}{2\sigma^{2}}\left\{-2\operatorname{tr}(\mathbf{R}\mathbf{B}\boldsymbol{\Psi}\Xi)+\operatorname{tr}(\Xi^{t}\boldsymbol{\Psi}^{t}\mathbf{B}^{t}\mathbf{B}\boldsymbol{\Psi}\Xi)\right\}-\frac{1}{2}\operatorname{tr}\{\mathbf{H}\boldsymbol{\Psi}^{t}\mathbf{P}_{\alpha}\boldsymbol{\Psi}\}\right]\times\mathbbm{1}(\boldsymbol{\Psi}\in\mathcal{V}_{K,Q})$
$\propto\exp\left\{\frac{1}{\sigma^{2}}\operatorname{tr}(\mathbf{R}\mathbf{B}\boldsymbol{\Psi}\Xi)-\frac{1}{2\sigma^{2}}\operatorname{tr}(\Xi^{t}\boldsymbol{\Psi}^{t}\mathbf{B}^{t}\mathbf{B}\boldsymbol{\Psi}\Xi)-\frac{1}{2}\operatorname{tr}(\mathbf{H}\boldsymbol{\Psi}^{t}\mathbf{P}_{\alpha}\boldsymbol{\Psi})\right\}\times\mathbbm{1}(\boldsymbol{\Psi}\in\mathcal{V}_{K,Q})$
$\propto\exp\left\{\frac{1}{\sigma^{2}}\operatorname{tr}(\Xi\mathbf{R}\mathbf{B}\boldsymbol{\Psi})-\frac{1}{2\sigma^{2}}\operatorname{tr}(\Xi\Xi^{t}\boldsymbol{\Psi}^{t}\mathbf{B}^{t}\mathbf{B}\boldsymbol{\Psi})-\frac{1}{2}\operatorname{tr}(\mathbf{H}\boldsymbol{\Psi}^{t}\mathbf{P}_{\alpha}\boldsymbol{\Psi})\right\}\times\mathbbm{1}(\boldsymbol{\Psi}\in\mathcal{V}_{K,Q})$
$\propto\exp\left\{\operatorname{tr}\left(\frac{\Xi\mathbf{R}\mathbf{B}\boldsymbol{\Psi}}{\sigma^{2}}-\frac{\Xi\Xi^{t}\boldsymbol{\Psi}^{t}\mathbf{B}^{t}\mathbf{B}\boldsymbol{\Psi}}{2\sigma^{2}}-\frac{\mathbf{H}\boldsymbol{\Psi}^{t}\mathbf{P}_{\alpha}\boldsymbol{\Psi}}{2}\right)\right\}\times\mathbbm{1}(\boldsymbol{\Psi}\in\mathcal{V}_{K,Q})$
The final line above is the form found in Result 2, which is not of a known distributional family.
S4 Simulation results
S4.1 Background
@@
S4.2 Fixed effects comparison
We compare the fixed effects estimates of $\mu(t)$ produced by FAST to those from GFSR and VMP. POLAR is not included, as it subtracts off the column means from teh data matrix in pre-processing rather than estimating the mean as part of the model. We compare ISE of the posterior estimate and aggregate coverage of equal-tailed $95\%$ credible intervals, each calculated using the same procedures outlined for the FPCs in Section 4.
@@
Figure S3: Kernel smoother of coverage probabilities of 9595% credible intervals of the true mean $\mu(t)$ for FAST and GFSR. Row indicates simulation scenario, S1 then S2. Distribution means: horizontal solid lines; nominal 9595% level: horizontal dotted lines.
@@
Figure S4 visualizes the distribution of score MSE over simulations. FAST is has consistently similar or lower MSE when compared to the existing methods. GFSR has much larger MSE for those scores associated with $\phi_{1}(t)$ in S1. The order of this MSE is close to $140^{2}=19600$, supporting our previous assertion that the mean $\mu(t)$ is being projected onto $\phi_{1}(t)$ and $\phi_{3}(t)$ by this implementation.
Figure S5: Kernel smoother of coverage probabilities of 9595% credible intervals of the true scores, $\xi_{ik}^{b}$, for FAST, GFSR, and VMP. First row: S1; second row: S2. Each column: corresponding eigenfunction. Distribution means: horizontal solid lines; nominal 9595% level: horizontal dotted lines.
Figure S5 displays the kernel smooth of the vectors of coverage proportions for FAST and relevant comparator methods by FPC (columns) and simulation scenario (rows). FAST is the only implementation which produces consistently close to nominal coverage for all FPCs under both scenarios.
S4.4 Multilevel extension
@@
$\phi_{k}^{(1)}(t)$ $=\{\sqrt{2}\sin(2\pi t),\sqrt{2}\cos(2\pi t),\sqrt{2}\sin(4\pi t),\sqrt{2}\cos(4\pi t)\};\quad\lambda^{(1)}_{k}=0.5^{k-1}$
$\phi_{l}^{(2)}(t)$ $=\{1,\sqrt{3}(2t-1),\sqrt{5}(6t^{2}-6t+1),\sqrt{7}(20t^{3}-30t^{2}+12t-1)\};\quad\lambda^{(2)}_{l}=0.5^{l-1}$
@@
Figure S7: Kernel smoother of coverage probabilities of 9595% credible intervals of the true FPCs for FAST and GFSR from the multilevel simulations. Row indicates hierarchical level, while columns correspond to FPCs. The horizontal dotted lines indicate the nominal 9595% level.
@@
Figure S8: Kernel smoother of coverage probabilities of 9595% credible intervals of the true scores, $\xi_{ik}^{b},\zeta_{ijk}^{b}$, for FAST and GFSR from the multilevel simulations. Row indicates hierarchical level, while columns correspond to FPCs. The horizontal dotted lines indicate the nominal 9595% level.
S5 Simulation sensitivity analyses
@@
S5.1 Chosen number of FPCs $K$
@@
Figure S9 indicates that the value of $K$ does not appear to have a substantial impact on the accuracy of FPC estimates, outside of not estimating FPCs when they are present. When $K=2$, the first two FPCs are still well estimated, and all 3 true FPCs are recovered with qualitatively similar accuracy whenever $K\geq 3$.
@@
Figure S10 shows that $K=2$ produces non-optimal inference for $\phi_{1}(t),\phi_{2}(t)$, with inferences for the first 3 FPCs unaffected when $K\geq 3$.
S5.2 Chosen spline dimension $Q$
@@
S5.3 Chosen penalty parameter $\alpha$
We finally vary penalty hyperparameter $\alpha$, which dictates what proportion of the smoothing penalty used by FAST will be allocated to absolute value rather than second-order variation. We vary between $\alpha=0.01$ and $\alpha=0.3$, keeping $\alpha<0.5$ to ensure the majority of the penalty is focused on the ”wiggliness” which should be the core of any smoothing penalty. We evaluate FAST under each $\alpha$ value using ISE and coverage of the FPCs as described in Section 4.
@@
S5.4 Timing by $Q,K$
Given that it appears that estimation accuracy and inference are consistent when $K$ and $Q$ are sufficiently large, we evaluate the effect of increasing these quantities on the computation time of FAST. For this experiment, we focus on S1, the more computationally complex simulation scenario based upon CGM data. For each combination of $K\in\{3,4,5\}$ and $Q\in\{20,30,40\}$, we timed FAST on the same personal laptop described in Section 4 (2023 MacBook Pro with Apple M2 Max@@3.49 GHz and 32GB of memory).
@@
From Table S1, we find that FAST does not require appreciable additional computation time when $Q,K$ are increased.
S6 Additional CGM Analyses/Details
S6.1 Variability explained - Single Level
@@
S6.2 Bayesian FPCA of randomly-chosen CGM within diets
@@
Figure S16: Bayesian FPCA results for the first three PCs (each column corresponds to one FPC) for each of the four diets (each row corresponds to one diet). X-axis: time from the start of the meal. Black curves: posterior mean; red curves: three samples from the posterior of the PCs; shaded areas: pointwise 9595% credible intervals.
@@
Figure S17: Eigenvalue and percent-variability estimates from the Bayesian FPCA models fit to each of the four diets (each line and color corresponds to one diet). X-axis: eigenfunction corresponding to eigenvalue/percent variance explained. All estimates are presented with their corresponding 9595% credible intervals.
As for the aggregate data, we fit each model using $2000$ iterations with the first $1000$ discarded as burn-in. Employing the routine from Section 2.5 to assess convergence of the scores and FPCs, we found final Gelman-Rubin statistics were <1.05<1.05.